# One Stone, Four Birds: A Comprehensive Solution for QA System Using Supervised Contrastive Learning

### v1

@@ VI-E Result on task 4: continual learning
Finally, we present the classification results of $ENC_{new}$ after the incremental learning in Table X. Overall, due to the excellent clustering performance on Task 3, this time our method further widened the accuracy gap with FT on all datasets. The concerning forgetting phenomenon also did not appear, as there was only a slight decrease on IND classification accuracy after the continual learning. This proves that our training strategy of using both pseudo-labeled IND and OOD texts is effective.
In addition to the continued high accuracy, we would like to highlight another training advantage of SCL over CE fine-tuning. In FT re-training, the last fully-connected layer had to be replaced and reinitialized to accommodate the new intents, requiring at least 4-5 epochs to obtain desirable results in our experiments. On the other hand, since SCL focuses on the feature optimization and label judgment is based on the nearest neighbor algorithm, it directly used the new data with extended labels for model training, offering straightforward scalability and efficiency.
In conclusion, our method demonstrated high standards of practicality and efficiency across all tasks, making it feasible to train and maintain a robust QA system quickly and easily.
VI-F Embedding space visualization
@@
Finally, to evaluate the representation optimization effect, we conducted a t-SNE embedding space visualization on BANKING Test I set. The results are shown in Figure 4. The top row displays the embedding arrangement over all intents (77 categories), while the bottom row focuses only on the newly discovered (previous OOD, 20 categories) data.
@@
From the visualization result on overall data distribution, it can be observed that both SCL and FT performed reasonably well, although FT mixed embeddings from different intents more frequently. On the other hand, MTP-KNNCL produced relatively larger clusters, with some clusters very close to each other, even starting to “stick” together. We consider this may be due to the training effect of KNN-ConL in MTP-KNNCL, which continuously pulled nearby similar embeddings together, resulting in a side effect on cluster aggregation.
@@
The advantage of our proposed method is more obvious in the visualization on newly discovered intents. Without accessing OOD texts, our method still separated most of the OOD classes, with only two class clusters mixed. MTP-KNNCL also worked relatively well, but its dispersion effect of different intents was inferior to ours. Finally, FT produced a flatter embedding distribution with more severe confusion between intents, leading to the lowest clustering scores among the three methods.
@@
To conclude, despite using less than half the time to the state-of-the-art method MTP-KNNCL and requiring fewer tuning steps, our method still produced a splendid underlying embedding space for clustering and other tasks, which is valuable in practical applications.
VII Conclusion
In this paper, we addressed the functionality and efficiency problems of current QA system, and proposed our solution to solve them and build a comprehensive system. We first defined four key tasks to complement the system’s functionality shortcomings, and then applied SCL as the unified representation learning method with minimal additional tuning to achieve the task targets. The experiment results demonstrated that although being simple and straightforward, the proposed method still achieved leading results in each task.
As this complete solution is first proposed and is still a prototype, there are still many situations needed to be considered for real deployment in the future. For example, there may also be some noisy inputs appearing, which are hard to be categorized into any intents. In such cases, developing a filtering algorithm before the intent discovery process might be necessary. In addition, to show efficiency, this time the downstream methods were all chosen to be as simple as possible, while more advanced methods could potentially yield even better results when combined with SCL.
@@ Appendix B Metric calculation details
Here we show the detailed calculation process of the used metrics.
@@ B-B For Task 3: NMI, ARI, ACC
Appendix C Abbreviations
Table XI shows the abbreviations that have appeared in this paper and their complete names.
TABLE XI: Abbreviations and full names
Abbreviations Full name
QA question answering (system)
RepL representation learning (feature learning)
ConL contrastive learning
MeL metric learning
UConL unsupervised contrastive learning
SCL supervised contrastive learning
IND in-domain (texts)
OOD out-of-domain (texts)
GDA Gaussian discriminant analysis
LOF local outlier factor
GPT Generative Pre-Trained Transformer
BERT Bidirectional Encoder Representations from Transformers
CE cross-entropy (loss)
FT fine-tuning
MLM masked language modeling
NN nearest neighbor
@@ Appendix D Learning and evaluation procedure (Algorithm form)
9:         Save model of this epoch, and record the threshold $\lambda$ when TPR=90%
10:     end if
11:end while
12:        -Task 1-
13:Input $\bm{x}_{\text{test I}}$ (IND part) to ENC to obtain the representations:
14:    $\bm{h}_{\text{test I}}=ENC(\bm{x}_{\text{test-1}})$
15:The label is determined as the label of the nearest IND cluster centroid:
16:    $\hat{y}_{\text{test I}}=\arg\min_{c\in\mathcal{C}}MDist(\bm{h}_{\text{test I}},\bar{\bm{h}}_{c})$
17:Calculate Micro, Macro F1 values
18:        -Task 2-
19:Input $\bm{x}_{\text{test-1}}$ (IND, OOD part) to ENC to obtain the representations and calculate the detection scores:
20:    $sco(\bm{x}_{\text{test I}})={\underset{c\in{\mathcal{C}}}{min}}\ MDist(\bm{h}_{\text{test I}},\bar{\bm{h}}_{c})$
21:Calculate FPR, TPR results on various thresholds $\lambda$, and finally AUROC, AUPR, FPR90 values
22:      -Task 3 preparation-
23:Determine $\lambda_{filter}$ when TPR=90% on $\bm{x}_{\text{val}}$
24:Divide $\bm{x}_{\text{test-1}}$ into $\bm{x}_{\text{test-1, IND}}$ and $\bm{x}_{\text{test-1, OOD}}$ two parts by comparing detection $sco$ and $\lambda_{filter}$
25:        -Task 3-
26:Apply KMeans on $\bm{x}_{\text{test-1, OOD}}$ and give new pseudo labels to $\bm{x}_{\text{test-1, OOD}}$ as $\hat{y}_{\text{test-1, OOD}}\in\mathcal{C}_{kn^{\prime}}$
27:Calculate NMI, ARI, ACC metrics
28:        -Task 4-
29:Give pseudo labels to $\bm{x}_{\text{test-1, IND}}$ using the initial ENC as $\hat{y}_{\text{test-1, IND}}\in\mathcal{C}_{kn}$
@@
31:Test on $\bm{x}_{\text{test II}}$ and calculate Micro, Macro F1 values
32:END

### v2

@@ VI-D Result on task 4: continual learning
VII Additional studies and discussions
VII-A Embedding space visualization
@@
To verify the representation optimization effect, we conducted a t-SNE embedding space visualization on BANKING Test I set, which has a moderate number of examples and categories to view results easily. The results are shown in Figure 4. The top row displays the embedding arrangement over all intents, while the bottom row focuses only on the newly discovered (i.e., previous OOD intents) data.
@@
From the visualization results on overall data distribution (Figure 4 a-c), it can be observed that both SCL and FT performed reasonably well, but the embedding space generated by FT showed more mixed embeddings from different intents. This proves that our proposed method has a better cluster aggregation effect over traditional FT. MTP-KNNCL produced relatively larger clusters in the space with some clusters very close to each other, but the overall distribution situation is slightly better than FT. We consider this may be due to the optimization style of KNN-ConL part used in MTP-KNNCL, which continuously pulled nearby similar embeddings together, resulting in this side effect on cluster aggregation.
@@
From the visualization results of newly discovered intents (Figure 4 d-f), the advantage of proposed method is more obvious. Without accessing OOD texts, SCL was still able to separate most OOD classes, with only two class clusters apparently mixed. MTP-KNNCL also worked relatively well, but the clusters were gathered closer, which showed an inferior labeling dispersion effect to our method. Finally, FT produced a much flatter embedding distribution with more severe confusion between intents. These visualization results correspond exactly to the results in Task 3.
@@
To conclude, the visualization result proved again that despite using less than half the time to the state-of-the-art method MTP-KNNCL and requiring fewer tuning steps, our method still produced a splendid underlying embedding space for clustering and other tasks, which is valuable in practical applications.
VII-B Computational efficiency analysis
Computational efficiency refers to the time or resources required for model training. In this analysis, we use the total computation amount required for a complete training procedure as an indicator of computational efficiency, which can be represented as:
$Eff.\propto Amount\;=\;Epoch\ast(O(model)+O(loss))$ (15)
where O(model) denotes the complexity of model’s forward and back propagation, and O(loss) represents the computation complexity of the loss function.
Since all methods in the experiment used the same base encoder, the O(model) part of them can be considered to be approximately equal [^2] Although GPT-2-small and BERT-base differ, their parameter sizes are close, so we consider their O(model) to be nearly equal.. In addition, given that O(model) is usually significantly larger than O(loss) (The analysis of O(loss) of each method and O(model) can be found in Appendix C), eq. 15 can be simplified to:
$\begin{split}&Eff.\propto Amount\propto Epoch\ast O(model)\\ &\rightarrow Eff.\propto Amount\propto Epoch\end{split}$ (16)
This simplification indicates that the computational efficiency, represented by computation amount, is primarily related to the number of training epochs. Reviewing the results from the tasks (Tables IV-VI), it is evident that the proposed SCL framework consistently converged in fewer epochs and required less training time, and thus demonstrating its advantage on computational efficiency.
Among the baselines, PTO-Sup and Layer-LOF also appeared to be efficient in terms of epoch count, but their training times were significantly longer. We consider the reason as PTO-Sup required complete sentence generation and loss calculation for each word, leading to substantial asynchronous processing delays. Layer-LOF, while not inherently inefficient, used adversarial attacks to oversample the data, which multiplied the training load. These factors inflated their total computation and resulted in suboptimal efficiency.
VII-C Potential challenges, limitations and future solutions
While our method demonstrated strong performance across tasks, as a prototype, it still presents several challenges and limitations needed to be considered for practical deployment.
Background-shift in data: In T-2 and T-3, despite better results than FT, the supervised nature of SCL can still misalign OOD texts with background-shifts, drawing them closer to IND clusters. To address this, future work could include pre-evaluating background-shift intensity and minimizing pre-training epochs to balance IND intent classification and OOD detection accuracy. Incorporating unsupervised training targets may also help.
Noisy inputs: The filtering and re-training strategy between T-3 and T-4 could struggle with noisy inputs that resist categorization, complicating KMeans clustering. The solution could be using advanced clustering methods like HDBSCAN, which can identify outliers without requiring a fixed cluster number. Alternatively, a filtering algorithm could be implemented to iteratively remove distant or scattered data points, though it risks accumulating noises over time.
Scalability on larger datasets: Task 4 results showed good scalability with the addition of up to 38 new intents. However, as category numbers increase, the model’s ability to distinguish between highly similar classes may diminish. When this phenomenon is clearly observed, splitting the model into sub-models based on broader categories may be beneficial, and we plan to explore optimal splitting strategies in future work.
VIII Conclusion
Traditional classification-retrieval-based QA systems primarily focus on categorizing known queries, which is insufficient to address the complexities of real-world scenarios. Moreover, the efficiency of the default training method is often suboptimal. To address these issues, we introduced three new tasks — out-of-domain detection, new intent discovery, and continual learning — to enhance the system’s functional capabilities. We then applied SCL as a unified representation learning method with minimal additional tuning to efficiently achieve the objectives of these tasks. Experimental results demonstrated that, despite its simplicity, the proposed approach consistently outperformed state-of-the-art methods across all tasks. These improvements, integrated within a natural workflow, significantly boost system adaptability and robustness, giving a meaningful impact on the future development of real-world QA systems.
While the proposed method is effective, it remains a prototype with room for further enhancement. Besides the challenges addressed in Section VII-C, future work could focus on integrating more advanced downstream methods to harness the full potential of SCL. For example, a larger-parameter encoder could provide a stronger representation base for learning, while more sophisticated analysis algorithms like energy-based methods should leverage SCL’s tighter cluster formations for enhanced OOD detection effects. In addition, applying advanced techniques like DeepCluster [59] for new intent discovery and Elastic Weight Consolidation [60] for continual learning could also further strengthen the system’s performance. Finally, the proposed SCL framework can also be extended to other QA tasks seamlessly, such as providing top-K confidence results for ambiguous classifications or visualizing sub-intents through embedding space analysis.
Ethical and privacy statement
In this study, the datasets used are all public-available and do not contain privacy-sensitive or confidential information. The neural networks used (i.e., BERT-series) are also less likely to inadvertently disclose sensitive information compared to generative models like GPTs.
Acknowledgment
The authors would like to thank Seita Shimada and Yoshiaki Matsukawa, Rakuten Card Co. Ltd., for their helpful comments on this study.
@@ Appendix A KMeans details
TABLE IX: Abbreviations and full names
Abbreviations Full name
QA question answering (system)
RepL representation learning (feature learning)
ConL contrastive learning
MeL metric learning
UConL unsupervised contrastive learning
SCL supervised contrastive learning
IND in-domain (texts)
OOD out-of-domain (texts)
GDA Gaussian discriminant analysis
LOF local outlier factor
GPT Generative Pre-Trained Transformer
BERT Bidirectional Encoder Representations
from Transformers
CE cross-entropy (loss)
FT fine-tuning
MLM masked language modeling
NN nearest neighbor
@@ Appendix B Metric calculation details
Here we show the detailed calculation processes of the used metrics.
@@ B-B For Task 3: NMI, ARI, ACC
Appendix C Computational complexity analysis
The computational complexity analysis of each method is shown in Table VIII according to the original papers.
Here, $N$ is the number of total data, $N_{b}$ is the number of data in one mini-batch and $N_{layer}$ is the number of layers in the network, where for bert-base it is 12. $H$ is the vector dimension of the hidden layer, for bert-base it is 768. $C$ is the number of IND categories. $L$ stands for the average input sentence length, while $V$ is the number of words in dictionary (specially for PTO-Sup). Finally, $K_{adv}$ denotes the number of adversarial examples generated (specially for Layer-LOF), $K_{knn}$ is the number of neighbors used in contrastive learning (specially for MTP-KNNCL) and $K_{proto}$ is the number of assumed prototype categories (specially for DPN).
@@ Appendix D Learning and evaluation procedure (Algorithm form)
Appendix E Abbreviations
Table IX shows the abbreviations that have appeared in this paper and their complete names.
@@ Appendix E Abbreviations
9:         $AUROC_{best-val}=AUROC_{val}$
10:         Save model of this epoch, and record the threshold $\lambda$ when TPR=90%
11:     end if
12:     $Epoch=Epoch+1$
13:end while
14:        -Task 1-
15:Input $\bm{x}_{\text{test I}}$ (IND part) to ENC to obtain the text representations:
16:    $\bm{h}_{\text{test I}}=ENC(\bm{x}_{\text{test I}})$
17:The label is determined as the label of the nearest IND cluster centroid:
18:    $\hat{y}_{\text{test I}}=\arg\min_{c\in\mathcal{C}_{kn}}MDist(\bm{h}_{\text{test I}},\bar{\bm{h}}_{c})$
19:Calculate Micro, Macro F1 values
20:        -Task 2-
21:Input $\bm{x}_{\text{test I}}$ (IND, OOD part) to ENC to obtain the representations and calculate the detection scores:
22:    $sco(\bm{x}_{\text{test I}})={\underset{c\in{\mathcal{C}}_{kn}}{min}}\ MDist(\bm{h}_{\text{test I}},\bar{\bm{h}}_{c})$
23:Calculate FPR, TPR results on various thresholds $\lambda$, and finally AUROC, AUPR, FPR90 values
24:      -Task 3 preparation-
25:Determine $\lambda_{filter}$ when OOD detection TPR=90% on $\bm{x}_{\text{val}}$
26:Divide $\bm{x}_{\text{test I}}$ into $\bm{x}_{\text{test I, `IND'}}$ and $\bm{x}_{\text{test I, `OOD'}}$ two parts by comparing detection $sco$ and $\lambda_{filter}$
27:        -Task 3-
28:Apply KMeans on $\bm{x}_{\text{test I, `OOD'}}$ and give new pseudo labels to $\bm{x}_{\text{test I, `OOD'}}$ as $\hat{y}_{\text{test I, `OOD'}}\in\mathcal{C}_{kn^{\prime}}$
29:Calculate NMI, ARI, ACC metrics
30:        -Task 4-
31:Give pseudo labels to $\bm{x}_{\text{test I, `IND'}}$ using the initial ENC as $\hat{y}_{\text{test I, `IND'}}\in\mathcal{C}_{kn}$
@@
33:Test on $\bm{x}_{\text{test II}}$ and calculate Micro, Macro F1 values
34:END